ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 146, 2005

Aljadeff, E., D. Haile and M. Natapov Projective bases of division algebras and groups of central type

317

Champetier, C. and V. Guirardel Limit groups as limits of free groups

1

Colombo, J. and P. Koshlukov Identities with involution for the matrix algebra of order two in characteristic p

337

Demers, M. F. Markov extensions for dynamical systems with holes: An application to expanding maps of the interval

189

Friedlander, L. Genericity of simple eigenvalues for a metric graph 149

Guirardel, V. See C. Champetier

Haile, D. See E. Aljadeff

Jiao, Z. The quasitriangular structures for a class of T -smash product Hopf algebras

125

Kang, M.-C. Some group actions on K(x_{1}, x_{2}, x_{3})  77

Koshlukov, P. See J. Colombo

Lackenby, M. Expanders, rank and graphs of groups  357

Lehavi, D. Any smooth plane quartic can be reconstructed from its bitangents

371

Leibman, A. Convergence of multiple ergodic averages along polynomials of several variables

303

Li, H. and L. Vrancken A basic inequality and new characterization of Whitney spheres in a complex space form

223

Liao, L. W. and Z. Ye A class of second order differential equations 281
McCutcheon, R. FVIP systems and multiple recurrence  157

Moretó, A. Large orbits of p-groups on characters and applications to character degrees

243

Natapov, M. See E. Aljadeff

Putnam, I. F. Lifting factor maps to resolving maps  253
Ransford, T. Eigenvalues and power growth  93

Vrancken, L. See H. Li

Ye, Z. See L. W. Liao Zalesskii, P. A. Profinite surface groups and the congruence kernel

of arithmetic lattices in SL_{2}(R)

111